# Strong discrete Morse theory

### v1

@@ 3. Strong discrete Morse theory
To establish that $M$ is acyclic, consider any pair $(\sigma,\{a\}\cup\sigma)\in M$. For any $\tau\succ\sigma$, we claim that $(\tau\smallsetminus\{a^{\prime}\},\tau)\notin M$ for any $a^{\prime}\neq a$. Since $\tau\succ\sigma$, the simplex $\tau$ contains $v_{\sigma}$. If $v_{\sigma}$ also satisfies $v_{\sigma}=\arg\max\{f(v)\colon v\in\tau\text{ is a descending dominated vertex}\}$, then $(\tau,\{a\}\cup\tau)\in M$, as $a$ dominates $v_{\sigma}$. Alternatively, if there exists another dominated vertex $w\in\tau$ such that $w>v_{\sigma}$, then $(\tau,\{a_{w}\}\cup\tau)\in M$, where $a_{w}$ dominates $w$. In either case, $(\tau\smallsetminus\{a^{\prime}\},\tau)\notin M$ for any $a^{\prime}\neq a$, ensuring that $M$ is acyclic. This completes the proof. ∎
@@ 4. The construction of the internal core
The critical poset can be constructed algorithmically, as detailed in Algorithm 1.
Algorithm 1 The critical poset
2:Critical poset ${\mathrm{Crit}}_{M}(K)$
3:$P\leftarrow$ elements of $\mathcal{X}(K)$
4:$R\leftarrow$ cover relations of $\mathcal{X}(K)$
5:Define a map ${\mathrm{QuotientClass}}:P\to\text{new elements}$
6:for all $(\sigma,\tau)\in M$ do
7:     $P\leftarrow P\setminus\{\sigma,\tau\}$
9:     $P\leftarrow P\cup\{[\sigma\sim\tau]\}$
10:     ${\mathrm{QuotientClass}}[\sigma]\leftarrow[\sigma\sim\tau]$
11:     ${\mathrm{QuotientClass}}[\tau]\leftarrow[\sigma\sim\tau]$
12:end for
14:     ${\mathrm{QuotientClass}}[\sigma]\leftarrow\sigma$
15:end for
16:$R^{\prime}\leftarrow\emptyset$
17:for all $(\sigma,\tau)\in R$ do
20:     end if
21:end for
24:return ${\mathrm{Crit}}_{M}(K)$
@@ 5. Algorithmic simplification of simplicial complexes
To compute the minimal strong internal core, we iteratively apply a greedy algorithm that performs randomly chosen internal strong collapses whenever a dominated vertex is available. If no dominated vertex exists, we randomly select a vertex to be removed. Algorithm 2 formalizes this procedure. As part of this procedure, we construct the acyclic matching associated to the internal strong collapses as stated in Theorem 3.9. Notice that Algorithm 2 can also be used to study $(N+1)$-deformations, as the strong internal core $(N+1)$-deforms to the original complex. A similar procedure can be applied for the randomized computation of a minimal weak core of a simplicial complex.
Remark 5.1.
Algorithm 2 A random strong internal core
${\mathcal{X}}(K)$: the face poset of a simplicial complex $K$
The critical poset of a random strong internal core of $K$
function StrongMorseReduction($X,C,M$)
     if $X=\emptyset$ then
         return $(X,C,M)$
     end if
     for $v\in S$ do
                  $M\leftarrow M\cup\{(\sigma,\{a\}\cup\sigma)\}$
              end for
              $X\leftarrow X\setminus\{\tau\in X\colon v\leq\tau\}$
              return StrongMorseReduction($X,C,M$)
         end if
     end for
     $C\leftarrow C\cup\{\sigma\in X\colon v\leq\sigma\}$
     $X\leftarrow X\setminus\{\sigma\in X\colon v\leq\sigma\}$
     return StrongMorseReduction($X,C,M$)
end function
procedure ComputeInternalStrongCore(${\mathcal{X}}(K)$)
     return CriticalPoset(${\mathcal{X}}(K),C,M$)
end procedure
@@ Acknowledgments
The author thanks Robert Green for his enthusiasm for posets and for pointing out the earlier work of Nicholas Scoville on Strong Morse Theory during his visit to Oxford. This work was inspired by fruitful conversations with him at the Mathematical Institute. The author is also grateful to Vidit Nanda for engaging with early versions of this project and for suggesting the connection with the localization of categories.

### v2

@@ 3. Strong discrete Morse theory
To establish that $M$ is acyclic, consider any pair $(\sigma,\{a\}\cup\sigma)\in M$. For any $\tau\succ\sigma$, we claim that $(\tau\smallsetminus\{a^{\prime}\},\tau)\notin M$ for any $a^{\prime}\neq a$. Since $\tau\succ\sigma$, the simplex $\tau$ contains $v_{\sigma}$. If $v_{\sigma}$ also satisfies $v_{\sigma}=\arg\max\{g(v)\colon v\in\tau\text{ is a descending dominated vertex}\}$, then $(\tau,\{a\}\cup\tau)\in M$, as $a$ dominates $v_{\sigma}$. Alternatively, if there exists another dominated vertex $w\in\tau$ such that $w>v_{\sigma}$, then $(\tau,\{a_{w}\}\cup\tau)\in M$, where $a_{w}$ dominates $w$. In either case, $(\tau\smallsetminus\{a^{\prime}\},\tau)\notin M$ for any $a^{\prime}\neq a$, ensuring that $M$ is acyclic. This completes the proof. ∎
Remark 3.10 (Connection to Bestvina–Brady Morse theory).
Let $K$ be a finite simplicial complex. A Bestvina–Brady Morse function $h\colon|K|\to\mathbb{R}$ is given by an injective function $g\colon V(K)\to\mathbb{R}$ that extends linearly over simplices. In this setting, the critical points of $h$ are precisely the vertices of $K$, and the changes in homotopy type of sublevel sets occur at critical levels by attaching cones over the descending links of the corresponding critical vertices. In our framework, we also consider a vertex function $g$, but classify each vertex as either strong critical or descending dominated, depending on whether its descending link forms a simplicial cone. This distinction not only describes the local homotopy behaviour (descending dominated vertices have coned descending links and, hence, do not change the homotopy type of sublevel complexes), but crucially allows a global, purely combinatorial reconstruction of the reduced complex ${\mathrm{core}}_{g}(K)$.
Forman’s discrete Morse theory fits within the Bestvina–Brady framework, as observed by Zaremsky [38], and can also be interpreted through the lens of strong discrete Morse theory. Specifically, any injective [^3] Any discrete Morse function can be replaced by an injective one with the same set of critical simplices and matched pairs; see the beginning of the proof of Theorem 3.3 in [19]. discrete Morse function $f\colon K\to\mathbb{R}$ induces a Bestvina-Brady Morse function $g\colon V(K^{\prime})\to\mathbb{R}$ such that critical simplices of $f$ correspond to vertices in $K^{\prime}$ whose descending links are barycentric subdivisions of the boundaries of those simplices — and hence homeomorphic to spheres. Each matched pair $(\sigma,\tau)$, with $\sigma\prec\tau$ and $f(\sigma)>f(\tau)$, satisfies that $\sigma$ is descending dominated by $\tau$, and ${\mathrm{lk}}^{\downarrow}(\tau,K^{\prime})$ is a topological cone on the unique vertex $v\in\tau\setminus\sigma$ (though not necessarily a simplicial one). This suggests a natural extension of our theory to cases where strong critical vertices have descending links that are topological cones.
@@ 4. The construction of the internal core
For details on the algorithmic construction of the critical poset, see Appendix A, Algorithm 1.
@@ 5. Algorithmic simplification of simplicial complexes
To compute the minimal strong internal core, we iteratively apply a greedy algorithm that performs randomly chosen internal strong collapses whenever a dominated vertex is available. If no dominated vertex exists, we randomly select a vertex to be removed (see Appendix A, Algorithm 2 for further details). As part of this procedure, we construct the acyclic matching associated to the internal strong collapses as stated in Theorem 3.9. Notice that this algorithm can also be applied to study $(N+1)$-deformations, as the strong internal core $(N+1)$-deforms to the original complex. A similar procedure can be applied for the randomized computation of a minimal weak core of a simplicial complex.
@@
Appendix A Algorithms
This appendix contains the pseudocode of the main algorithms used in the construction and computation of the critical poset associated with a strong internal core. An implementation in SageMath [16] can be found at https://github.com/ximenafernandez/Strong-Morse-Theory.
Algorithm 1 The critical poset
2:Critical poset ${\mathrm{Crit}}_{M}(K)$
3:$P\leftarrow$ elements of $\mathcal{X}(K)$
4:$R\leftarrow$ cover relations of $\mathcal{X}(K)$
5:Define a map ${\mathrm{QuotientClass}}:P\to\text{new elements}$
6:for all $(\sigma,\tau)\in M$ do
7:     $P\leftarrow P\setminus\{\sigma,\tau\}$
9:     $P\leftarrow P\cup\{[\sigma\sim\tau]\}$
10:     ${\mathrm{QuotientClass}}[\sigma]\leftarrow[\sigma\sim\tau]$
11:     ${\mathrm{QuotientClass}}[\tau]\leftarrow[\sigma\sim\tau]$
12:end for
14:     ${\mathrm{QuotientClass}}[\sigma]\leftarrow\sigma$
15:end for
16:$R^{\prime}\leftarrow\emptyset$
17:for all $(\sigma,\tau)\in R$ do
20:     end if
21:end for
24:return ${\mathrm{Crit}}_{M}(K)$
Algorithm 2 A random strong internal core
${\mathcal{X}}(K)$: the face poset of a simplicial complex $K$
The critical poset of a random strong internal core of $K$
function StrongMorseReduction($X,C,M$)
     if $X=\emptyset$ then
         return $(X,C,M)$
     end if
     for $v\in S$ do
                  $M\leftarrow M\cup\{(\sigma,\{a\}\cup\sigma)\}$
              end for
              $X\leftarrow X\setminus\{\tau\in X\colon v\leq\tau\}$
              return StrongMorseReduction($X,C,M$)
         end if
     end for
     $C\leftarrow C\cup\{\sigma\in X\colon v\leq\sigma\}$
     $X\leftarrow X\setminus\{\sigma\in X\colon v\leq\sigma\}$
     return StrongMorseReduction($X,C,M$)
end function
procedure ComputeInternalStrongCore(${\mathcal{X}}(K)$)
     return CriticalPoset(${\mathcal{X}}(K),C,M$)
end procedure
Remark A.1.
@@ Acknowledgments
The author thanks Robert Green for his enthusiasm for posets and for pointing out the earlier work of Nicholas Scoville on strong Morse theory during his visit to Oxford. This work was inspired by fruitful conversations with him at the Mathematical Institute. The author is also grateful to Vidit Nanda for engaging with early versions of this project and for suggesting the connection with the localization of categories, and to Gabriel Minian for useful comments on the connection with Bestvina–Brady Morse theory, and to Eugenio Borghini for valuable discussions. This work was partially supported by Leverhulme Trust Research Project Grant RPG-2023-144 and the UK Centre for Topological Data Analysis EPSRC grant EP/R018472/1.